# Multivariate generalized Pareto distributions along extreme directions

### v2

@@ 1 Introduction
In the literature, tail dependence is described mathematically through various concepts, such as the exponent measure, the angular measure, and the stable tail dependence function, among others. When the focus is on modeling multivariate threshold exceedances, the mgp distribution is a natural candidate (Rootzén and Tajvidi, 2006). In the univariate case, peaks-over-threshold modeling with the generalized Pareto distribution is well-established (Davison and Smith, 1990), while in the multivariate set-up, statistical modeling is quite recent (Rootzén et al., 2018; Kiriliouk et al., 2018; Ho and Dombry, 2019; Engelke and Hitz, 2020; Hentschel et al., 2022).
Current theory for mgp distributions does not cover extreme directions other than $J=\{1,\ldots,d\}$, a gap we aim to fill. From their generators, we derive expressions for mgp densities with respect to an appropriate dominating measure.
@@ 2 Background
and let $\pi_{J}:\mathbb{A}_{J}\to\mathbb{R}^{\lvert J\rvert}$ be the projection $\boldsymbol{x}\mapsto\boldsymbol{x}_{J}$
@@ 2.1 Multivariate extreme value distributions
Many functions and measures that describe this dependence are discussed in the literature (see, e.g., Beirlant et al., 2004, Chapter 8).
@@
where $G^{*}$ is an mev distribution with unit-Fréchet margins. For $\boldsymbol{\xi}_{i}$ as in (2.2), if $\boldsymbol{X}_{i}=\left({X_{i1},\ldots,X_{id}}\right)$ is defined by then we have $\max_{i=1,\ldots,n}\boldsymbol{X}_{i}/n\rightsquigarrow G^{*}$ as $n\to\infty$.
@@
The measure $\mu$ is finite on Borel sets whose topological closure does not contain the origin, i.e., is bounded away from the origin. Moreover, $\mu\left({\mathbb{E}\setminus\left[\boldsymbol{0},\boldsymbol{x}\right]}\right)=\infty$ as soon as $x_{j}=0$ for some $j\in D$. The exponent measure $\mu$ in the sense that $\mu(c\,\cdot\,)=c^{-1}\mu(\,\cdot\,)$ for all $c>0$. Moreover, for any $\mu$-continuous Borel set $B\subset\mathbb{E}$ bounded away from the origin, we have
@@ 2.2 Multivariate generalized Pareto distributions
where $\boldsymbol{\zeta}$ is the vector of lower endpoints of $G_{1},\ldots,G_{d}$ and $\vee$ denotes the component-wise maximum. We write $H=\operatorname{GP}(G)$. As in Rootzén et al. (2018), we consider the standardized case where $G$ has Gumbel margins, i.e., $\gamma_{j}=\mu_{j}=0$ and $\alpha_{j}=1$ for all $j\in D$. The distribution of an mgp random vector $\boldsymbol{Y}\sim H$ is then determined by the stdf $\ell$.
@@
Let $\boldsymbol{S}$ be a random vector in $[-\infty,0]^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{S})=0]=1$, $\operatorname{\mathsf{P}}[S_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{S_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{S_{d}}]$. Then $\boldsymbol{S}$ is said to be the spectral random vector associated to the mgp $H$ if
@@
with $E=\max(\boldsymbol{Y})$ a unit exponential random variable independent of the spectral random vector $\boldsymbol{S}=\boldsymbol{Y}-\max(\boldsymbol{Y})$.
Conversely, any random vector $\boldsymbol{S}$ with the above properties generates an mgp $H$.
@@
Let $\boldsymbol{T}$ be a random vector in $\left[-\infty,\infty\right)^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{T})>-\infty]=1$, $\operatorname{\mathsf{P}}[T_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{T_{1}-\max(\boldsymbol{T})}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{T_{d}-\max(\boldsymbol{T})}]$. The random vector $\boldsymbol{T}$ is called a $\boldsymbol{T}$-generator of $H$ , if we have the equality in distribution
@@
Let $\boldsymbol{U}$ be a random vector on $[-\infty,\infty)^{d}$ such that $\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{U_{d}}]$ and $0<\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]<\infty$. The random vector $\boldsymbol{U}$ is said to be an $\boldsymbol{U}$-generator of $H$ if
@@ 3 Extreme directions
and assume the multivariate regular variation condition $n\operatorname{\mathsf{P}}\left[{\boldsymbol{X}\in n\,\cdot\,}\right]\to\mu(\,\cdot\,)$ in the sense of (2.4) holds. We are interested in scenarios where all variables $X_{j}$ for $j$ in some subset $J$ of $D$ are large while the other variables are not; a precise meaning will be given below. Such a group of variables $J$ will be called an extreme direction, defined formally in Section 3.1. We will provide various characterizations of such extreme directions: by the angular measure, by the asymptotic probability of $\boldsymbol{X}$ to belong to certain sets called thickened rectangles and cones, and finally by the mgp distribution. Next, in Section 3.2, we express the conditional probability density function of an mgp random vector given a certain extreme direction in terms of other densities that are easier to compute.
@@ 3.1 Definition and characterizations
Recall the exponent measure $\mu$ of $G^{*}$ as defined in (2.3) and $\mathbb{E}=[\boldsymbol{0},\boldsymbol{\infty})\setminus\left\{{\boldsymbol{0}}\right\}$. In this section, let $J\subseteq D$ be a fixed non-empty set and $\mathbb{E}_{J}:=\left\{{\boldsymbol{x}\in\mathbb{E}:\ x_{j}>0\text{ iff }j\in J}\right\}$. The collection $\left\{{\mathbb{E}_{J}:\;\emptyset\neq J\subseteq D}\right\}$ forms a partition of $\mathbb{E}$. The following definition was introduced in Simpson et al. (2020) in order to identify extreme directions of the random vector $\boldsymbol{X}$ defined as in (LABEL:equa:standardization).
@@
The tail dependence coefficient $\chi_{J}$ in (LABEL:eq:chiJ) satisfies
@@ 4.1 Definition and properties
In case $Z_{1}^{(k)}=\ldots=Z_{d}^{(k)}$ almost surely for all $k\in R$, the distribution of $\boldsymbol{M}$ is max-linear with unit-Fréchet margins . The mixture model (4.1) can thus be seen as a smoothed version of the max-linear factor model in the sense that complete dependence within each of the factors $\boldsymbol{Z}^{(1)},\ldots,\boldsymbol{Z}^{(r)}$ is relaxed to a possibly weaker dependence structure.
@@
The random vector $\boldsymbol{U}$ in (4.7) will be called a mixture generator with coefficient matrix $A$, generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ and mass vector $\boldsymbol{m}$.
@@ 5.2.3 Simulation example
associated with the mixture logistic and Hüsler–Reiss models in panels (a) and (b), respectively. The Box–Cox type transformation in (5.7) ensures that the vector of lower endpoints of $\boldsymbol{Z}$ is $\boldsymbol{\eta}=(-4,-4,-4)$ and facilitates visualization.

### v3

@@ 1 Introduction
In the literature, tail dependence is described mathematically through various concepts, such as the exponent measure, the angular measure, and the stable tail dependence function, among others. When the focus is on modeling multivariate threshold exceedances, the mgp distribution is a natural candidate (Rootzén and Tajvidi, 2006). In the univariate case, peaks-over-threshold modeling with the generalized Pareto distribution is well-established (Davison and Smith, 1990), while in the multivariate set-up, statistical modeling is quite recent (Rootzén et al., 2018; Kiriliouk et al., 2018; Ho and Dombry, 2019; Engelke and Hitz, 2020; Hentschel et al., 2022). Asymptotically, the class of mgp distributions represents proper multivariate probability distributions on a region where at least one variable is large. Moreover, parametric mgp models can be constructed quite easily via the generators proposed in Kiriliouk et al. (2018).
Current theory for mgp distributions does not cover extreme directions other than $J=\{1,\ldots,d\}$, a gap we aim to fill. From their generators, we derive expressions for mgp densities with respect to an appropriate dominating measure. Furthermore, we propose a stochastic construction based on smoothed max-linear models, wherein the strong dependence within factors inherent in max-linear models is substituted with a more lenient dependency framework. Our construction is designed to accommodate all mgp models, regardless of their extreme directions.
@@
Our contribution differs from those above from several perspectives. First, our focus is on extreme directions in the context of multivariate threshold exceedances. Second, our aim is not to learn extreme directions, but rather to introduce parametric mgp models capable of generating extreme directions other than $\left\{{1,\ldots,d}\right\}$. For multivariate extreme-value distributions, the max-linear factor model (Einmahl et al., 2012) and the asymmetric logistic model (Tawn, 1990) already allow for this possibility. In the recent literature, only Simpson et al. (2020) and Chiapino et al. (2020) propose mixture models for multivariate extreme-value distributions. The power of our model construction resides in the fact that it is capable of representing mgp distributions with general extreme directions. Moreover, for specific parametric families, our model is straightforward to simulate from.
@@ 2 Background
Let $d$ be a positive integer and write $D=\left\{{1,\ldots,d}\right\}$. Throughout, bold symbols will refer to multivariate quantities. Let $\boldsymbol{0}$ and $\boldsymbol{1}$ denote vectors of zeroes and ones, respectively, of a dimension clear from the context. A $d$-dimensional vector $\boldsymbol{x}$ will be written as $\boldsymbol{x}=(x_{1},\ldots,x_{d})$ and for a non-empty set $J\subseteq\left\{{1,\ldots,d}\right\}$, write $\boldsymbol{x}_{J}=(x_{j})_{j\in J}$. Mathematical operations on vectors such as addition, multiplication and comparison are considered component-wise apart from $\boldsymbol{a}\nleqslant\boldsymbol{b}$ which denotes $a_{j}>b_{j}$ for at least one $j\in D$.
@@
and let $\pi_{J}:\mathbb{A}_{J}\to\mathbb{R}^{\lvert J\rvert}$ be the projection $\boldsymbol{x}\mapsto\boldsymbol{x}_{J}$, where $\lvert J\rvert$ denotes the number of elements of the set $J$.
@@ 2.1 Multivariate extreme value distributions
For $d\geqslant 2$, the family of $d$-dimensional mev distributions no longer constitutes a finite-dimensional parametric family, in contrast to the univariate case. The reason is the dependence structure between the components of $G$, or equivalently, the extremal dependence between the components of $\boldsymbol{\xi}$. Many functions and measures that describe this dependence are discussed in the literature (see, e.g., Beirlant et al., 2004, Chapter 8). In this paper, three such objects will appear on the forefront: the exponent measure, the angular measure (de Haan and Resnick, 1977) and the stable tail dependence function (Drees and Huang, 1998).
To concentrate on the dependence structure of $G$, it is convenient to standardize its margins. The choice of common scale is a matter of convention and convenience. In this section, we will opt for the unit-Fréchet distribution, i.e., $\alpha_{j}=\gamma_{j}=\mu_{j}=1$ for all $j\in D$, in line with the theory as originally exposed in (de Haan and Resnick, 1977). In Section 2.2, however, we will choose the Gumbel distribution instead ($\alpha_{j}=1$ and $\gamma_{j}=\mu_{j}=0$ for all $j\in D$), in order to link up with the additive model formulation for the mgp in Rootzén et al. (2018). Switching from one standardization to another is an easy and sometimes illuminating exercise.
@@
where $G^{*}$ is an mev distribution with unit-Fréchet margins. For $\boldsymbol{\xi}_{i}$ as in (2.2), if $\boldsymbol{X}_{i}=\left({X_{i1},\ldots,X_{id}}\right)$ is defined by
$X_{ij}=-1/\ln F_{j}\left({\xi_{ij}}\right),\qquad i\in\mathbb{N},\;j\in D,$ (2.3)
then we have $\max_{i=1,\ldots,n}\boldsymbol{X}_{i}/n\rightsquigarrow G^{*}$ as $n\to\infty$.
@@
The measure $\mu$ is finite on Borel sets whose topological closure does not contain the origin, i.e., is bounded away from the origin. Moreover, $\mu\left({\mathbb{E}\setminus\left[\boldsymbol{0},\boldsymbol{x}\right]}\right)=\infty$ as soon as $x_{j}=0$ for some $j\in D$. The exponent measure $\mu$ is $(-1)$-homogeneous in the sense that $\mu(c\,\cdot\,)=c^{-1}\mu(\,\cdot\,)$ for all $c>0$. Moreover, for any $\mu$-continuous Borel set $B\subset\mathbb{E}$ bounded away from the origin, we have
@@ 2.2 Multivariate generalized Pareto distributions
where $\boldsymbol{\zeta}$ is the vector of lower endpoints of $G_{1},\ldots,G_{d}$ and $\vee$ denotes the component-wise maximum. The conditioning event $\boldsymbol{\xi}\nleqslant\boldsymbol{b}_{n}$ is to be read as $\exists j\in D:\xi_{j}>b_{nj}$, that is, at least one variable within $\boldsymbol{\xi}$ exceeds a high threshold. We write $H=\operatorname{GP}(G)$. As in Rootzén et al. (2018), we consider the standardized case where $G$ has Gumbel margins, i.e., $\gamma_{j}=\mu_{j}=0$ and $\alpha_{j}=1$ for all $j\in D$. The distribution of an mgp random vector $\boldsymbol{Y}\sim H$ is then determined by the stdf $\ell$.
@@
Let $\boldsymbol{S}$ be a random vector in $[-\infty,0]^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{S})=0]=1$, $\operatorname{\mathsf{P}}[S_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{S_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{S_{d}}]$. Then $\boldsymbol{S}$ is said to be the spectral random vector (Rootzén et al., 2018, Theorem 6) associated to the mgp $H$ if
@@
with $E=\max(\boldsymbol{Y})$ a unit exponential random variable independent of the spectral random vector $\boldsymbol{S}=\boldsymbol{Y}-\max(\boldsymbol{Y})$; see Rootzén et al. (2018, Theorem 7).
The distribution of the random vector $\boldsymbol{S}$ can thus be obtained directly from the mgp $H$. Conversely, any random vector $\boldsymbol{S}$ with the above properties generates an mgp $H$. Still, the constraint $\max(\boldsymbol{S})=0$ almost surely is not satisfied by common families of distributions, which raises the question how to construct appropriate $\boldsymbol{S}$. The next representation resolves this issue, thereby facilitating the construction of an mgp.
@@
Let $\boldsymbol{T}$ be a random vector in $\left[-\infty,\infty\right)^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{T})>-\infty]=1$, $\operatorname{\mathsf{P}}[T_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{T_{1}-\max(\boldsymbol{T})}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{T_{d}-\max(\boldsymbol{T})}]$. The random vector $\boldsymbol{T}$ is called a $\boldsymbol{T}$-generator of $H$ (Rootzén et al., 2018, Proposition 8), if we have the equality in distribution
@@
Any random vector $\boldsymbol{T}$ with the above properties yields an mgp random vector $\boldsymbol{Y}$ via (2.13). In comparison to $\boldsymbol{S}$, it has one constraint less to satisfy; the construction in (2.12) guarantees that $\max(\boldsymbol{S})$ vanishes almost surely. On the other hand, the $\boldsymbol{T}$-generator of an mgp distribution $H$ is not unique: adding the same random variable to all components of $\boldsymbol{T}$ does not change the distribution of $\boldsymbol{T}-\max(\boldsymbol{T})$ and produces the same mgp $H$.
A drawback of both representations $\boldsymbol{S}$ and $\boldsymbol{T}$ is that they are not stable with respect to marginalization. If $\boldsymbol{Y}\sim H$ is a $d$-dimensional mgp random vector and if $J\subset D$ is non-empty, the conditional distribution of $\boldsymbol{Y}_{J}$ given $\max\boldsymbol{Y}_{J}>0$ is a $|J|$-dimensional mgp, say $H_{|J}$ (which is different from $H_{J}$, the unconditional distribution of $\boldsymbol{Y}_{J}$, which is in general not a mgp). However, if $\boldsymbol{S}$ is the spectral random vector of $H$, then $\boldsymbol{S}_{J}$ is not the spectral random vector of $H_{|J}$. The following representation resolves this issue, at the price of a change of measure.
@@
Let $\boldsymbol{U}$ be a random vector on $[-\infty,\infty)^{d}$ such that $\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{U_{d}}]$ and $0<\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]<\infty$. The random vector $\boldsymbol{U}$ is said to be an $\boldsymbol{U}$-generator of $H$ (Rootzén et al., 2018, Proposition 9) if
@@
This constraint still does not make $\boldsymbol{U}$ identifiable from $H$: if the random variable $V$ satisfies $\operatorname{\mathsf{E}}\left[{\mathrm{e}^{V}}\right]=1$ and is independent of $\boldsymbol{U}$, then the random vector $(U_{1}+V,\ldots,U_{d}+V)$ yields another $\boldsymbol{U}$-generator for the same mgp $H$.
If $\boldsymbol{U}$ is a generator of $\boldsymbol{Y}\sim H$, then, for non-empty $J\subset D$, the random vector $\boldsymbol{U}_{J}$ is a generator of $\boldsymbol{Y}_{J}\mid\max(\boldsymbol{Y}_{J})>0$, which thus follows a $|J|$-variate mgp (Rootzén et al., 2018, Proposition 18). In Proposition 5.2 below, we establish a link between $\boldsymbol{U}$ and the extremal functions in Dombry et al. (2016). Equation (2.14) represents the stdf $\ell$ as a $D$-norm with generator $\boldsymbol{U}$ in the sense of Falk (2019). In Kiriliouk et al. (2018, Section 7), common parametric families of multivariate extreme value models are described in terms of their $\boldsymbol{U}$-generator.
@@
Note that a spectral random vector $\boldsymbol{S}$ is also a $\boldsymbol{T}$-generator, since $\max(\boldsymbol{S})=0$ and thus $\operatorname{GP}_{\boldsymbol{T}}(\mathcal{L}(\boldsymbol{S}))=\operatorname{GP}_{\boldsymbol{S}}(\mathcal{L}(\boldsymbol{S}))$. On the other hand, the representation via $\boldsymbol{U}$ is of another nature: if $\boldsymbol{V}$ is a random vector whose distribution satisfies both the constraints to be a $\boldsymbol{T}$-generator and a $\boldsymbol{U}$-generator, then the mgp distributions $\operatorname{GP}_{\boldsymbol{T}}(\mathcal{L}(\boldsymbol{V}))$ and $\operatorname{GP}_{\boldsymbol{U}}(\mathcal{L}(\boldsymbol{V}))$ are different in general; see Kiriliouk et al. (2018, Section 7) for examples.
@@ 3 Extreme directions
Let the random vector $\boldsymbol{X}=(X_{j})_{j\in D}$ be defined on the basis of the probability integral transform applied to the components of a random vector $\boldsymbol{\xi}$ as in (2.3), and assume the multivariate regular variation condition $n\operatorname{\mathsf{P}}\left[{\boldsymbol{X}\in n\,\cdot\,}\right]\to\mu(\,\cdot\,)$ in the sense of (2.5) holds. We are interested in scenarios where all variables $X_{j}$ for $j$ in some subset $J$ of $D$ are large while the other variables are not; a precise meaning will be given below. Such a group of variables $J$ will be called an extreme direction, defined formally in Section 3.1. We will provide various characterizations of such extreme directions: by the angular measure, by the asymptotic probability of $\boldsymbol{X}$ to belong to certain sets called thickened rectangles and cones, and finally by the mgp distribution. Next, in Section 3.2, we express the conditional probability density function of an mgp random vector given a certain extreme direction in terms of other densities that are easier to compute.
Upper tail dependence coefficients (Coles et al., 1999; Schlather and Tawn, 2002)
$\chi_{J}=\lim_{q\to\infty}q\operatorname{\mathsf{P}}\left[{\forall j\in J:X_{j}>q}\right],\qquad\emptyset\neq J\subseteq D,$ (3.1)
are well-known summary measures of extremal dependence. However, we would like to emphasize that they are not sufficient to characterize extreme directions. In fact, even though the equality $\chi_{J}=0$ implies that the variables $X_{j}$ for $j\in J$ cannot be large simultaneously, the inequality $\chi_{J}>0$ does not exclude the possibility of extreme directions properly contained in $J$. For instance, for a pair $J=\left\{{j_{1},j_{2}}\right\}$, the inequality $\chi_{j_{1}j_{2}}>0$ implies that $X_{j_{1}}$ and $X_{j_{2}}$ can be large simultaneously but does not rule out the possibility that $X_{j_{1}}$ is large while $X_{j_{2}}$ is not.
@@ 3.1 Definition and characterizations
Recall the exponent measure $\mu$ of $G^{*}$ as defined in (2.4) and $\mathbb{E}=[\boldsymbol{0},\boldsymbol{\infty})\setminus\left\{{\boldsymbol{0}}\right\}$. In this section, let $J\subseteq D$ be a fixed non-empty set and $\mathbb{E}_{J}:=\left\{{\boldsymbol{x}\in\mathbb{E}:\ x_{j}>0\text{ iff }j\in J}\right\}$. The collection $\left\{{\mathbb{E}_{J}:\;\emptyset\neq J\subseteq D}\right\}$ forms a partition of $\mathbb{E}$. The following definition was introduced in Simpson et al. (2020) in order to identify extreme directions of the random vector $\boldsymbol{X}$ defined as in (2.3).
@@
The tail dependence coefficient $\chi_{J}$ in (3.1) satisfies
@@ 4.1 Definition and properties
In case $Z_{1}^{(k)}=\ldots=Z_{d}^{(k)}$ almost surely for all $k\in R$, the distribution of $\boldsymbol{M}$ is max-linear with unit-Fréchet margins and we are back to the factor model in Example 3.7. The mixture model (4.1) can thus be seen as a smoothed version of the max-linear factor model in the sense that complete dependence within each of the factors $\boldsymbol{Z}^{(1)},\ldots,\boldsymbol{Z}^{(r)}$ is relaxed to a possibly weaker dependence structure. Under the conditions of Proposition 4.11 below, the resulting mixture model has a density with respect to the measure $\upsilon$ in Eq. (3.9). We hope this property will facilitate statistical inference through likelihood-based methods.
@@
The random vector $\boldsymbol{U}$ in (4.7) will be called a mixture generator with coefficient matrix $A$, generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ and mass vector $\boldsymbol{m}$. Different choices of the vector $\boldsymbol{m}$ yield different $\boldsymbol{U}$ generators, all of which however induce the same mgp distribution $H$, namely, the one associated to the mixture model. Among all possible choices for $\boldsymbol{m}$, the obvious one is $m_{k}=1/r$ for all $k\in\left\{{1,\ldots,r}\right\}$, but the formula allows for other possibilities. Further, note that $m_{k}$ is not equal to the probability that the mgp vector $\boldsymbol{Y}\sim H$ lies in extreme direction $J_{k}$: instead, we have $\operatorname{\mathsf{P}}[Y_{j}>-\infty\text{ iff }j\in J_{k}]=w_{k}$, a quantity which is proportional to $\ell_{J_{k}}^{(k)}((a_{jk})_{j\in J_{k}})$, as stated in Proposition 5.1 below.
@@ 5.2.3 Simulation example
associated with the mixture logistic and Hüsler–Reiss models in panels (a) and (b), respectively. The Box–Cox type transformation in (5.7) ensures that the vector of lower endpoints of $\boldsymbol{Z}$ is $\boldsymbol{\eta}=(-4,-4,-4)$ and facilitates visualization. By Proposition 4.5, the matrix $A$ produces a mixture model with three distinct extreme directions $J\subset D$, namely $\left\{{1,2,3}\right\}$, $\left\{{2,3}\right\}$, and $\left\{{3}\right\}$. This is visible in Figure 2 by a concentration of points on the three corresponding subfaces $\mathbb{A}^{-\boldsymbol{4}}_{J}=\left\{{\boldsymbol{x}\in[-4,\infty)^{3}:\ x_{j}>-4\text{ iff }j\in J}\right\}$.
@@
To assess Algorithm 1, Table 1 compares the empirical proportions of points in the three sub-faces $\mathbb{A}_{J}^{-4}$ in a sample of size $1000$ with the corresponding true probabilities in Eq. (5.1) for the logistic and Hüsler–Reiss mixture models.